A stretched string has a mass per unit length of 3.86 g/cm and a tension of 25.2 N. A sinusoidal wave on this string has an amplitude of 0.137 mm and a frequency of 156 Hz and is traveling in the negative direction of an x axis. If the wave equation is of the form
y(x,t) = ym sin(kx + ωt),
what are (a) ym, (b) k, and (c) ω, and (d) the correct choice of sign in front of ω?
Answer #1

(a) ym = amplitude = .137 x 10^-3 m

(c) w = 2*pi*f
w = 2 x 3.14 x 15
  
w = 979.68

(b) k = w / v

v = (T/m)^1/2 = (25.2 / .386)^1/2 = 8.07

k = 979.68 / 8.07

k = 121.39

(d) It has to be "+"
   Because you follow a wave by fixing the value of the phase, and then seeing how x changes at t changes.